3. [-12 Points] DETAILS LARLINALG8 7.2.007. For the matrix A, find (if possible) a nonsingular matrix P such that p-1AP is di1. [0/2 Points] DETAILS PREVIOUS ANSWERS LARLINALG8 7.2.001. Consider the following. -11 40 A= -27 (a) Verify that A is diagoplease solve both
Answer #1

8 - 2 A -4 first we find eigen values of A. det IA-XI) 2o. 8- -2 zo 4 (8-))(1-2) – (8) 20. 8-80-ta2 -8=0 AM ²_9x=0 A(4-9)=0 afor X-9 (A ->I) x 20 8-9 - 2 (2 23][2] -4 1-9 - 2 al co 4 -8 y R₂ → R2-4R, - x - 2y =o. = ) E-3) (0,33 Y & = K₂ (let) 2y = -x2 PARS [:] [4}][: 112 8 -8 - 16 2 - 2 PAP [-4 9 - 4 - 4 +4 8 + 1 -18 -1] plap -18 +18 pt HAP 9 0 +72+9 5[4] [. [ Ś le to ;]Ques - -۱۱ -4 -5 A = [ - م [ ] - 2 + pla AP Diagonal Matrix with eigen values oot are the diagonal element. . 4 5 4 - (5) (s